### v1

# SlimCaching: Edge Caching of Mixture-of-Experts for Distributed Inference

Abstract
Mixture-of-Experts (MoE) models improve the scalability of large language models (LLMs) by activating only a small subset of relevant experts per input. However, the sheer number of expert networks in an MoE model introduces a significant storage burden for an edge device. To address this challenge, we consider a scenario where experts are dispersed within an edge network for distributed inference. Based on the popular Top-$K$ expert selection strategy, we formulate a latency minimization problem by optimizing expert caching on edge servers under storage constraints. When $K=1$, the problem reduces to a monotone submodular maximization problem with knapsack constraints, for which we design a greedy-based algorithm with a $(1-1/e)$-approximation guarantee. For the general case where $K\geq 1$, expert co-activation within the same MoE layer introduces non-submodularity, causing greedy methods to be ineffective. To tackle this issue, we propose a successive greedy decomposition method to decompose the original problem into a series of subproblems, with each being solved by a dynamic programming approach. Furthermore, we design an accelerated algorithm based on the max-convolution technique to obtain the approximate solution with a provable guarantee in polynomial time. Simulation results on various MoE models demonstrate that our method significantly reduces inference latency compared to existing baselines.
Index Terms:
Edge AI, large language models, mixture-of-experts, expert caching, edge inference.
I Introduction
Large language models (LLMs), such as GPT [1] and LLaMA [2], have delivered remarkable performance across various tasks, including many privacy-sensitive and real-time applications [3, 4, 5]. Driven by the growing demand for privacy preservation and low-latency responses, there is a pivotal trend toward deploying LLMs at the network edge [6, 7]. Reflecting this trend, leading technology companies such as Qualcomm, Huawei, and Apple have already integrated on-device LLMs into consumer-grade mobile devices [8, 9]. This movement toward edge deployment is expected to shape the future of LLMs and fundamentally transform the landscape of AI-powered mobile applications.
Despite these advances, LLMs at the network edge face inherent performance limitations compared with the cloud-based counterparts [10, 11]. This is because the capabilities of LLMs generally scale with model size and computing resources [12], making it challenging to achieve high performance under edge resource constraints. To overcome the scaling challenges, the Mixture-of-Experts (MoE) architecture has emerged as a predominant solution [13], which has been adopted in numerous state-of-the-art LLMs such as Phi-3.5-MoE, DeepSeek-V3, and Hunyuan-Large [14]. Specifically, MoE models replace each dense transformer block with a massive set of expert networks (e.g., feedforward networks (FFNs)), among which only a subset is activated for each input token, as visualized in Fig. 1. This paradigm allows LLMs to scale in size without a proportional increase in computation for each inference, thereby enabling an effective balance between model performance and computational efficiency – an essential requirement for deployment at the network edge.
However, while MoE architectures substantially reduce inference workload, they pose significant storage challenges for edge devices. For instance, the Switch Transformer architecture allows each MoE layer to host up to hundreds of experts, leading to an MoE model up to 65 times larger than a FLOPs-equivalent dense T5 model, from 446 MB to 29.4 GB [15]. Such an explosion in parameters makes them ill-suited for edge devices with limited storage capacity [16, 17], such as the iPhone 16 with only 128 GB basic storage. In practice, only a small fraction of experts can be cached locally on mobile devices, thereby constraining the scale of MoE models for on-device deployment.
Considering that the majority of parameters in MoE models are contributed by expert networks [^1] Considering the Switch Transformer model family, when the number of expert networks per MoE layer increases from 8 to 32, the proportion of the MoE model’s total data size occupied by the expert networks rises from 82.3% to 94.4%., a natural idea to address the above challenge is to distribute experts across edge networks. Specifically, we envision a scenario where users store only a subset of their preferred experts locally, whereas the wireless edge network caches the remaining experts. The wireless edge network, which can be a 6G mobile network, hosts a massive number of experts for various MoE-based LLMs on distributed edge servers (e.g., base stations). When desired experts are not locally available, users route their token embedding to nearby wireless edge servers with the corresponding experts for inference. This distributed inference framework for MoE models offer several salient advantages: 1) local data and final predictions remain on the user side with privacy benefits, akin to on-device inference, with only intermediate token embeddings uploaded to edge servers; 2) each user only stores minimal, “slim” version of the MoE models with their most frequently used experts, thus substantially reducing local storage and memory costs; 3) Compared with cloud-based MoE systems, edge caching of experts significantly reduce the response time and backbone traffic.
Figure 1: Visualization of activated experts in different MoE layers of the text part in MoE-LLaVA-Phi2-2.7B-4e with Top-2 strategy [18] under the Science Question Answering (SQA) dataset [19].
The paradigm above gives rise to a fundamental research question: Given the popularity of experts, how can they be optimally placed across distributed storage-constrained edge servers to enable latency-efficient MoE inference? At the first glance, the problem falls into traditional edge caching problems, which determine the placement of content on edge servers based on their content popularity. Nonetheless, key differences exist. In classical content caching or dense model placement, cached items are independently retrievable [20, 21, 22]. In contrast, the activated experts within the same MoE layer exhibit a strong correlation since the model activates $K$ experts per input (typically based on the Top-$K$ strategy). In particular, when multiple selected experts at the same layer are colocated on a single server, they require the identical input token embeddings, which can be transmitted only once. As a result, the overall layer latency is no longer a simple linear summation of individual expert latencies. The coupling effect among the selected experts introduces non-submodularity into the optimization problem, making it significantly more challenging. Consequently, existing delay-optimal content placement schemes, which are often based on submodular optimization and greedy procedures [23], are no longer effective for MoE placement in terms of theoretical approximation guarantees when $K>1$.
To address the challenging problem, in this paper, we design a distributed low-latency inference in MoE with expert caching (SlimCaching) framework under Top-$K$ strategy. Specifically, given the limited storage capacity of edge servers, we aim to determine optimal expert placement strategies that minimize the average inference latency across all users. The main contributions of this work are summarized as follows:
  • We define a novel expert caching problem tailored for distributed MoE inference. Considering a multi-edge system with cooperative caching between edge devices and edge servers, we formulate a combinatorial optimization problem to minimize the average inference latency subject to storage capacity constraints. We identify that the resulting problem is a monotone submodular maximization with multiple knapsack constraints when $K=1$, and becomes a monotone but non-submodular and non-supermodular maximization problem when $K>1$ due to expert dependencies.
  • For the case of $K=1$, we develop a greedy-based algorithm that achieves a $(1-1/e)$-approximation guarantee. For the more general setting with $K\geq 1$, we propose a successive greedy decomposition approach that reformulates the problem into a sequence of subproblems, each solvable via a dynamic programming (DP)-based algorithm. The proposed algorithm attains a $(1-\kappa_{g})/2$-approximate global solution in polynomial time, where $\kappa_{g}$ is the supermodular curvature. Furthermore, we propose an accelerated algorithm considering the uniform expert sizes within each MoE model.
  • We conduct extensive experiments on SQA and Visual Question Answering (VQA)-v2 datasets using diverse MoE models with varying Top-$K$ configurations. The results demonstrate that the proposed method consistently outperforms traditional greedy baselines in terms of both latency and computational efficiency.
The rest of this paper is organized as follows. We review the related works in Section II. Then, we provide the system models in Section III, including the expert activation procedure during MoE inference, expert caching model, and per-token inference latency. Based on the activation probability of experts, we formulate a latency-minimization optimization problem and characterize its structural property in Section IV. In Section V, we develop the algorithms for the special case $K=1$ and for the general case $K\geq 1$. Experimental results are presented in Section VI, followed by concluding remarks in Section VII.
II Related Works
The relevant research works on MoE and its edge deployment can be summarized as follows.
1) Expert selection: For MoE-based models, expert selection plays a central role in determining both inference latency and accuracy. The most popular method is Top-$K$ expert selection, where, for each input, the $K$ experts with the highest routing probabilities are chosen to process the data [24, 25]. There are also dynamic expert selection strategies, where a threshold is defined such that experts are selected if the cumulative routing probabilities exceed the threshold [26]. In this work, we adopt the widely used Top-$K$ expert selection to study the expert placement problem in edge networks.
2) In-memory expert caching: One of the main bottlenecks in MoE inference arises from the high memory overhead associated with storing a large number of expert networks. One usual practice is on-demand expert fetching, where inactive experts are kept in low-tier memory and only loaded into GPU memory when selected. However, since the router must first determine the Top-$K$ experts before initiating expert loading, the fetching process is fully serialized with computation, causing substantial latency. To address this issue, an expert prefetching mechanism is developed in [27, 28], where a transformer-based routing path predictor is trained offline to predict expert activation for each input token in a single pass. The predicted experts are then prefetched during the computation of previous layers. Furthermore, in [29, 30], by decoupling the gating router from the MoE architecture, tokens with similar activation patterns can be grouped into micro-batches, further reducing fetching latency.
3) In-network expert caching: Due to the huge sizes and computation burden, experts sometimes can be distributed over edge networks. By considering task relevance and wireless channel conditions, an energy-aware expert selection framework is developed, where experts are dispersed in wireless edge networks for distributed inference [31]. In this work, however, we focus on expert placement in edge networks based on the widely used Top-$K$ expert selection mechanism. To our best knowledge, the expert placement problem has not been investigated in the literature.
Expert placement is a key problem for in-network expert caching, i.e., which experts should be placed on which edge server. While caching has been extensively investigated in areas such as content delivery and model caching, expert caching in MoE models exhibits unique characteristics that render existing solutions inapplicable. In content caching, most prior work assumes independence among content items [32], and the objective is typically to maximize a submodular utility function subject to matroid constraints [23]. This setting enables the application of classical greedy algorithms with a well-known $(1-1/e)$-approximation guarantee. In cases where dependencies exist, such as in caching interdependent tasks modeled by directed acyclic graphs, content placement decisions are often made sequentially, with each task requiring a single function to be cached and executed [33, 34, 35, 36]. In model caching, a recent approach, TrimCaching [16, 37], exploits parameter sharing across AI models to improve storage efficiency. By treating model layers as cacheable units, the caching problem can be formulated as a supermodular maximization problem under multiple knapsack constraints. However, it assumes that a cached model is stored in its entirety, which is infeasible in the context of large-scale MoE models due to their considerable parameter sizes. Expert caching in MoE models poses fundamentally different challenges due to the Top-$K$ expert activation mechanism, which introduces two key difficulties: 1) Non-submodularity due to expert co-activation: When $K>1$, multiple experts must be activated simultaneously for each token, leading to strong co-activation dependencies among experts. The dependency causes the submodularity and supermodularity properties, which are commonly exploited in existing caching formulations, to fail to provide theoretical approximation guarantees in our expert placement problem. 2) Knapsack-type storage constraints: Expert caching is further complicated by the heterogeneity in expert sizes across different MoE models. Unlike the commonly assumed cardinality or matroid constraints, the expert caching problem introduces knapsack-type constraints that depend on the data size of cached experts. The above properties necessitate new problem formulations and optimization strategies for expert caching for distributed MoE inference.
III System Model
In this paper, we consider an MoE-based edge caching system with a set $\mathcal{N}=\left\{1,\ldots,N\right\}$ of edge servers and a set $\mathcal{U}=\left\{1,\ldots,U\right\}$ of users. The storage capacity of edge server $n\in\mathcal{N}$ is denoted by $Q_{n}$. Edge servers can communicate with each other directly via backhaul links and are also connected to the central cloud (expert model library) with a wired backhaul link. Each user connects to its nearest edge server and generates a task that requires an MoE architecture for inference, where $p_{u,m}$ denotes the probability of user $u$ requesting model $m$. There are $M$ requested MoE models, and the set of models is denoted by =$\mathcal{M}=\left\{1,\ldots,M\right\}$. For model $m$, we consider the Top-$K_{m}$ expert selection mechanism, where $K_{m}$ is the number of activated experts within an MoE layer.
III-A Preliminaries: Expert Activation During MoE Inference
We first briefly introduce the preliminaries of MoE inference. Fig. 2 shows the architecture of an MoE model. For MoE model $m\in\mathcal{M}$, the number of MoE layers is $L_{m}$, indexed by a set $\mathcal{L}_{m}=\left\{1,\ldots,L_{m}\right\}$. Each MoE layer consists of a gate network and a set of $E_{m}$ expert networks (i.e., FFNs), and the data size of a single expert is denoted by $b_{m}$. Thus, the total number of experts is given by $E=\sum_{m\in\mathcal{M}}L_{m}E_{m}$. We use $\mathcal{E}_{m}^{(\ell)}$ to denote the set of experts within model $m$’s MoE layer $\ell$.
Figure 2: Illustration of an MoE architecture with $E$ expert networks in each MoE layer.
We take model $m$ as an example to illustrate how expert(s) are activated at MoE layer $\ell$. Let ${\mathbf{g}}\left(\mathbf{s}\right)$ denote the token-to-expert affinity scores of an input $\mathbf{s}$, which represents the probability of the token selecting an expert for processing. Then, ${\mathbf{g}}\left(\mathbf{s}\right)$ can be expressed as
${\mathbf{g}}\left(\mathbf{s}\right)=\mathrm{Softmax}\left(\mathbf{W}_{r}\mathbf{s}^{\mathrm{T}}\right),$ (1)
where $\mathbf{W}_{r}$ is the router parameter. With Top-$K_{m}$ routing strategy, the experts with $K_{m}$ highest scores will be activated. Then, the probability of expert $i$ can be normalized as follows:
$\bar{{\mathbf{g}}}\left(\mathbf{s}\right)_{i}=\begin{cases}\frac{{\mathbf{g}}\left(\mathbf{s}\right)_{i}}{\sum_{i^{\prime}\in\mathrm{Top}K_{m}\left({\mathbf{g}}\left(\mathbf{s}\right)\right)}{\mathbf{g}}\left(\mathbf{s}\right)_{i^{\prime}}},&\text{ if }i\in{\mathrm{Top}}K_{m}\left({\mathbf{g}}\left(\mathbf{s}\right)\right),\\ 0,&\text{ otherwise. }\end{cases}$ (2)
Subsequently, the final output of the expert networks can be obtained by ${\mathbf{y}}\left(\mathbf{s}\right)=\sum_{i\in{\mathcal{E}}_{m}^{(\ell)}}{\bar{\mathbf{g}}\left(\mathbf{s}\right)_{i}y_{i}\left(\mathbf{s}\right)}$, where $y_{i}\left(\mathbf{s}\right)$ is the output result of the selected expert $i$.
Let $\mathcal{A}_{m}^{(\ell)}$ denote the collection of all possible expert index sets of size $K_{m}$ that can be activated in MoE layer $\ell$ of model $m$, i.e., $\mathcal{A}_{m}^{\left(\ell\right)}=\left\{S_{m}^{\left(\ell\right)}\subseteq\mathcal{E}_{m}^{\left(\ell\right)}\mid\left|S_{m}\right|=K_{m}\right\}$. For example, when $K_{m}=2$ and there are 3 experts in MoE layer $\ell$ of model $m$, we have $\mathcal{A}_{m}^{\left(\ell\right)}=\left\{\left\{1_{m}^{\left(\ell\right)},2_{m}^{\left(\ell\right)}\right\},\left\{1_{m}^{\left(\ell\right)},3_{m}^{\left(\ell\right)}\right\},\left\{2_{m}^{\left(\ell\right)},3_{m}^{\left(\ell\right)}\right\}\right\}$ and one possible value of $S_{m}^{\left(\ell\right)}$ is $\left\{1_{m}^{\left(\ell\right)},2_{m}^{\left(\ell\right)}\right\}$.
III-B Expert Caching Model
Given that expert networks account for the majority of MoE model’s parameters, we assume that all components of MoE models except the experts are deployed on users, as shown in Fig. 3. This strategy ensures raw input data and final predictions remain on the user side, thereby enhancing user privacy. Due to storage limitations, users can only store a limited number of experts in an MoE model. When the requested experts are not locally available, the user should send the token embeddings to edge servers with cached experts for inference. When the requested experts are not cached at any edge server, the user should send token embeddings to the cloud, which is assumed to cache all experts. After the requested experts complete processing, the nodes send the output results back to the user. Then, the processing of the next MoE layer begins.
Figure 3: A distributed wireless system for MoE inference, where users and edge servers have expert caching capabilities. (a) Illustration of local cache hit, edge cache hit, and edge cache miss. (b) Operations and protocol of SlimCaching within an MoE layer.
In this paper, we consider that experts on users are pre-determined according to user preference (e.g., their most frequently used experts), whereas the placement of experts in edge networks should be optimized according to user requests and network status. Let $\rho_{u,i_{m}^{\left(\ell\right)}}$ denote the expert caching state of the users. If the expert with index $i$ of MoE model $m$’s layer $\ell$ is pre-cached on user $u$, $\rho_{u,i_{m}^{\left(\ell\right)}}=1$ holds; otherwise, $\rho_{u,i_{m}^{\left(\ell\right)}}=0$. Let $x_{n,i_{m}^{\left(\ell\right)}}$ denote the caching decision variable for the edge servers, with $x_{n,i_{m}^{\left(\ell\right)}}=1$ indicating the expert with index $i$ of MoE model $m$’s layer $\ell$ is placed at edge server $n$ and $x_{n,i_{m}^{\left(\ell\right)}}=0$ otherwise.
III-C Per-token Inference Latency
Let $R_{u,n}^{\rm{UL}}$ denote the expected data rate from user $u$ to its associated (nearest) edge server $n_{u}$, which can be expressed as
$R_{u,n_{u}}^{\rm{UL}}=B_{u}\log_{2}\left(1+\frac{P_{u,n_{u}}G_{u,n_{u}}d_{u,n_{u}}^{-\alpha}}{N_{0}}\right),$ (3)
where $B_{u}$ is the bandwidth occupied by user $u$, $P_{u,n_{u}}$ is the transmit power from user $u$ to edge server $n_{u}$, $G_{u,n_{u}}$ is the uplink antenna-related factor, $d_{u,n_{u}}$ is the distance between user $u$ to edge server $n_{u}$, $\alpha$ is the path-loss coefficient, and $N_{0}$ is the noise power. Similarly, the downlink achievable rate from the edge server $n_{u}$ to user $u$ is given by
$R_{u,n_{u}}^{\rm{DL}}=B_{u}\log_{2}\left(1+\frac{P_{n_{u},u}G_{n_{u},u}d_{u,n_{u}}^{-\alpha}}{N_{0}}\right),$ (4)
where $P_{n_{u},u}$ is the transmit power from edge server $n_{u}$ to user $u$ and $G_{n_{u},u}$ is the downlink antenna-related factor. For the backhaul transmissions, let $R_{n,n^{\prime}}$ denote the transmission rate from edge server $n$ to $n^{\prime}$, and let $R_{n,\mathrm{C}}$ denote the transmission rate from edge server $n$ to the cloud.
When performing MoE inference, the output tensor preserves the same dimensions as the input tensor. Thus, when using MoE model $m$ for task inference, the uplink and downlink have identical data transmission sizes, i.e., the size of token embedding denoted by $D_{m}$. When using model $m$ for inference, given the uplink data rate in (3), the communication latency of transmitting a token embedding from user $u$ to its associated edge server $n_{u}$ is given by
$T_{u,n_{u},m}^{\rm{UL}}=\frac{D_{m}}{R_{u,n_{u}}^{\mathrm{UL}}}.$ (5)
Similarly, the communication latency of transmitting a token embedding via the downlink between user $u$ and edge server $n_{u}$, via backhaul link from edge server $n$ to $n^{\prime}$, and via backhaul link from edge server $n$ to the cloud can be expressed as $T_{u,n_{u},m}^{\rm{DL}}=\frac{D_{m}}{R_{u,n_{u}}^{\mathrm{DL}}}$, $T_{n,n^{\prime},m}=\frac{D_{m}}{R_{n,n^{\prime}}}$, and $T_{n,{\mathrm{C}},m}=\frac{D_{m}}{R_{n,{\mathrm{C}}}}$, respectively.
Let $\mu_{m}^{\mathrm{CP}}$ denote the computation workload (in FLOPs) of an expert in model $m$, which can be calculated according to [38]. Let $C_{u}$ be the computing capability (in FLOP/s) of user $u$, and the allocated computing capability for each expert denoted by ${\dot{C}}_{u}$ is given by ${\dot{C}}_{u}=C_{u}/E_{m}$. Then, the local computing latency of user $u$ utilizing the expert network of model $m$ can be expressed as
$T_{u,m}^{{\mathrm{CP}}}=\frac{\mu_{m}^{\mathrm{CP}}}{{\dot{C}}_{u}}=\frac{\mu_{m}^{\mathrm{CP}}E_{m}}{C_{u}}.$ (6)
Let ${\dot{C}}_{\mathrm{E}}$ and ${\dot{C}}_{\mathrm{C}}$ denote the allocated computing capability for each expert at the edge server and at the cloud, respectively. In the considered scenario, the users’ request traffic load is sufficiently low, such that computing resource contention among users is not considered. Thus, the computing latency of the edge server and the cloud utilizing the expert network of model $m$ can be given by $T_{\mathrm{E},m}^{{\mathrm{CP}}}=\frac{\mu_{m}^{\mathrm{CP}}}{{\dot{C}}_{\mathrm{E}}}$ and $T_{\mathrm{C},m}^{{\mathrm{CP}}}=\frac{\mu_{m}^{\mathrm{CP}}}{{\dot{C}}_{\mathrm{C}}}$, respectively.
Consider that the bottleneck of MoE inference in distributed wireless networks is the communication overhead of inter-nodes, and the delay of token routing to the cloud is much longer than that to edge networks [39]. Thus, the order in which user $u$’s token searches for expert processing is as follows: local user, associated (nearest) edge server $n_{u}$, other non-associated edge servers $\mathcal{N}\setminus n_{u}$, and finally the cloud. Fig. 4 shows four different cases of token routing when a user requests the expert pair $(i_{m}^{\ell},j_{m}^{\ell})$ under Top-2 strategy.
Figure 4: Different cases of token routing when $K=2$.
Consider an arbitrary user, say $u$, requires an expert (group) $S_{m}^{\left(\ell\right)}$ in the MoE layer $l$ of model $m$ for inference. Next, we will discuss the per-token inference latency when traversing this layer. The number of activated experts at different nodes can be calculated as follows.
  1. 1.
    At the local user $u$, there are $\beta_{u,S_{m}^{\left(\ell\right)}}=\mathop{\sum}\limits_{i_{m}^{\left(\ell\right)}\in S_{m}^{\left(\ell\right)}}\rho_{u,i_{m}^{\left(\ell\right)}}$ activated experts.
  2. 2.
    At the associated edge server $n_{u}$, there are $\beta_{u,n_{u},S_{m}^{\left(\ell\right)}}=\mathop{\sum}\limits_{i_{m}^{\left(\ell\right)}\in S_{m}^{\left(\ell\right)}}\left(1-\rho_{u,i_{m}^{\left(\ell\right)}}\right)x_{n_{u},i_{m}^{\left(\ell\right)}}$ activated experts.
  3. 3.
    At the cloud, there are $\beta_{u,\mathrm{C},S_{m}^{\left(\ell\right)}}=\sum_{i\in S_{m}^{\left(\ell\right)}}\left(1-\rho_{u,i_{m}^{\left(\ell\right)}}\right)\prod_{n\in\mathcal{N}}\left(1-x_{n,i_{m}^{\left(\ell\right)}}\right)$ activated experts.
  4. 4.
    At other edge servers excluding the associated edge server, there are $\beta_{u,S_{m}^{\left(\ell\right)}}^{\mathrm{OE}}=K_{m}-\beta_{u,S_{m}^{\left(\ell\right)}}-\beta_{u,n_{u},S_{m}^{\left(\ell\right)}}-\beta_{u,\mathrm{C},S_{m}^{\left(\ell\right)}}$ activated experts in total.
For the non-associated edge servers $n^{\prime}\in\mathcal{N}\setminus n_{u}$, let $\beta_{u,n^{\prime},i_{m}^{\left(\ell\right)}}\in\{0,1\}$ denote whether it provides processing with its cached expert $i_{m}^{\left(\ell\right)}$ for user $u$. Here, $\beta_{u,n^{\prime},i_{m}^{\left(\ell\right)}}\leq x_{n^{\prime},i_{m}^{\left(\ell\right)}}$ holds since edge server $n^{\prime}$ can provide inference only if it caches the requested expert. Then, the total number of utilized experts at the non-associated edge server $n^{\prime}$ is given by $\sum_{i_{m}^{\left(\ell\right)}\in S_{m}^{\left(\ell\right)}}\beta_{u,n^{\prime},i_{m}^{\left(\ell\right)}}$, and $\beta_{u,S_{m}^{\left(\ell\right)}}^{\mathrm{OE}}=\sum_{n^{\prime}\in\mathcal{N}\setminus n_{u}}\sum_{i_{m}^{\left(\ell\right)}\in S_{m}^{\left(\ell\right)}}\beta_{u,n^{\prime},i_{m}^{\left(\ell\right)}}$ holds. Here, $\beta_{u,n^{\prime},i_{m}^{\left(\ell\right)}}$ is determined in a way that leads to the lowest per-token inference latency, which can be easily obtained given expert placement decisions.
We assume that tokens are transmitted sequentially between the associated edge server $n_{u}$ and other edge servers. As a result, $T_{n_{u},S_{m}^{\left(\ell\right)}}^{\mathrm{OE}}$ is given by
$\begin{split}T_{n_{u},S_{m}^{\left(\ell\right)}}^{\mathrm{OE}}=\sum_{n^{\prime}\in\mathcal{N}\setminus n_{u}}&\left[{\mathbb{I}\left({{\beta_{u,n^{\prime},S_{m}^{\left(\ell\right)}}}>0}\right)\left({{T_{{n_{u}},n^{\prime},m}}+T_{{\mathrm{E}},m}^{{\mathrm{CP}}}}\right)}\right.\\ &\left.+{\beta_{u,n^{\prime},S_{m}^{\left(\ell\right)}}}{T_{n^{\prime},{n_{u}},m}}\right].\end{split}$ (7)
In a special case where $K_{m}=1$, the token only requires one expert in each MoE layer for processing. In this case, $S_{m}^{(\ell)}$ reduces to $i_{m}^{(\ell)}$ and $\beta_{u,n^{\prime},S_{m}^{(\ell)}}$ is given by
$\begin{split}\beta_{u,n^{\prime},S_{m}^{(\ell)}}&=\left(1-\rho_{u,S_{m}^{\left(\ell\right)}}\right)\left(1-x_{n_{u},S_{m}^{\left(\ell\right)}}\right)\\ &\cdot\mathop{\prod}\limits_{\begin{subarray}{c}n^{\prime\prime}\in\mathcal{N}\\ T_{n_{u},n^{\prime\prime},m}+T_{n^{\prime\prime},n_{u},m}\\ <T_{n_{u},n^{\prime},m}+T_{n^{\prime},n_{u},m}\end{subarray}}\left(1-x_{n^{\prime\prime},S_{m}^{\left(\ell\right)}}\right)x_{n^{\prime},S_{m}^{\left(\ell\right)}}.\end{split}$ (8)
That is, expert $S_{m}^{\left(\ell\right)}$ is cached at edge server $n^{\prime}$, and all other edge servers $n^{\prime\prime}\in\mathcal{N}$ that provide a lower round-trip transmission latency between $n_{u}$ and $n^{\prime\prime}$ (i.e., $T_{n_{u},n^{\prime\prime},m}+T_{n^{\prime\prime},n_{u},m}<T_{n_{u},n^{\prime},m}+T_{n^{\prime},n_{u},m}$) do not cache this expert. Therefore, the token from user $u$ is transmitted to edge server $n^{\prime}$ for processing.
Let us get back to the general case. Based on (7), the latency experienced by a token at the edge servers denoted by $T_{n_{u},S_{m}^{\left(\ell\right)}}^{\mathrm{E}}$ can be expressed as
$T_{n_{u},S_{m}^{\left(\ell\right)}}^{\mathrm{E}}=\mathbb{I}\left(\beta_{u,n_{u},S_{m}^{\left(\ell\right)}}>0\right)T_{\mathrm{E},m}^{\mathrm{CP}}+T_{n_{u},S_{m}^{\left(\ell\right)}}^{\mathrm{OE}}.$ (9)
In addition, the latency from the associated edge server $n_{u}$ to the cloud denoted by $T_{n_{u},S_{m}^{\left(\ell\right)}}^{\mathrm{C}}$ can be expressed as
$\begin{split}T_{n_{u},S_{m}^{\left(\ell\right)}}^{\mathrm{C}}&=\mathbb{I}\left({{\beta_{u,\mathrm{C},S_{m}^{\left(\ell\right)}}}>0}\right)\left({{T_{{n_{u}},\mathrm{C},m}}+T_{{\mathrm{C}},m}^{{\mathrm{CP}}}}\right)\\ &+{\beta_{u,\mathrm{C},S_{m}^{\left(\ell\right)}}}T_{\mathrm{C},{n_{u}},m}.\end{split}$ (10)
As a result, when the local experts cannot satisfy the inference requests, we can obtain the inference latency requiring outer nodes denoted by $T_{u,S_{m}^{\left(\ell\right)}}^{\mathrm{out}}$ as follows,
$\begin{split}T_{u,S_{m}^{\left(\ell\right)}}^{\mathrm{out}}&=T_{u,n_{u},m}^{\mathrm{UL}}+\left(K_{m}-\sum_{i_{m}^{\left(\ell\right)}\in S_{m}^{\left(\ell\right)}}\rho_{u,i_{m}^{\left(\ell\right)}}\right)T_{u,n_{u},m}^{\mathrm{DL}}\\ &+T_{n_{u},S_{m}^{\left(\ell\right)}}^{\mathrm{E}}+T_{n_{u},S_{m}^{\left(\ell\right)}}^{\mathrm{C}}.\end{split}$ (11)
Finally, the per-token inference latency when user $u$ requests expert(s) $S_{m}^{\left(\ell\right)}$ for inference can be expressed as
$T_{u,S_{m}^{\left(\ell\right)}}^{\mathrm{token}}=\prod_{i_{m}^{\left(\ell\right)}\in S_{m}^{\left(\ell\right)}}\rho_{u,i_{m}^{\left(\ell\right)}}T_{u}^{\rm{CP}}+\left(1-\prod_{i_{m}^{\left(\ell\right)}\in S_{m}^{\left(\ell\right)}}\rho_{u,i_{m}^{\left(\ell\right)}}\right)T_{u,S_{m}^{\left(\ell\right)}}^{\mathrm{out}}.$ (12)
III-D Activation Probability and Average Latency
As analyzed in Section I, when using different MoE models, each token will have a unique distribution of activation expert paths. Let $p_{u,S_{m}^{(\ell)}}$ denote the probability of user $u$ selecting expert(s) $S_{m}^{(\ell)}$ in model $m$’s MoE layer $\ell$ for processing, which can be measured by historical statistics.
Let ${\mathcal{S}}_{m,j}=\left\{S_{m,j}^{(1)},\dots,S_{m,j}^{(L_{m})}\right\}\in\mathcal{S}_{m}$ denote the $j$-th path when selecting model $m$ for inference, where $S_{m,j}^{(\ell)}\in{\mathcal{S}}_{m,j}$ is the activated expert(s) in the $\ell$-th MoE layer. The probability of each individual path ${\mathcal{S}}_{m,j}$ for user $u$, denoted by $p_{u,{\mathcal{S}}_{m,j}}$, can be expressed as
$\begin{split}p_{u,{\mathcal{S}}_{m,j}}&=p_{u}\left(S_{m,j}^{(1)},\dots,S_{m,j}^{(L_{m})}\right)\\ &=p_{u}\left(S_{m,j}^{(1)}\right)\prod_{\ell=2}^{L_{m}}p_{u}\left(S_{m,j}^{(\ell)}\mid S_{m,j}^{(1)},\ldots,S_{m,j}^{(\ell-1)}\right),\end{split}$ (13)
which is derived from the transition probability.
Let ${\bar{T}}_{u,m}$ denote the average per-token inference latency when user $u$ requires model $m$, which can be expressed as
$\begin{split}{\bar{T}}_{u,m}&=\sum_{{\mathcal{S}}_{m,j}\in\mathcal{S}_{m}}p_{u,{\mathcal{S}}_{m,j}}\sum_{S_{m,j}^{(\ell)}\in{\mathcal{S}}_{m,j}}T_{u,S_{m,j}^{(\ell)}}^{\mathrm{token}}\\ &=\sum_{\ell\in\mathcal{L}_{m}}\sum_{S_{m}^{(\ell)}\in{\mathcal{A}}_{m}^{(\ell)}}p_{u,S_{m}^{(\ell)}}T_{u,S_{m}^{(\ell)}}^{\mathrm{token}},\end{split}$ (14)
where the second equation is derived based on the characteristics of joint probability.
IV Problem Formulation and Characterization
In this section, we first formulate an optimization problem to minimize the average inference latency subject to the storage capacity constraints. Then, we identify and characterize the structural properties of the optimization problem under different $K$ values.
IV-A Problem Formulation
Define the latency reduction function as
$r_{u,S_{m}^{(\ell)}}\left(\mathbf{X}\right)\triangleq T_{u,S_{m}^{\left(\ell\right)}}^{\max}-T_{u,S_{m}^{(\ell)}}^{\mathrm{token}}\left({\mathbf{X}}\right),$ (15)
where
$T_{u,S_{m}^{\left(\ell\right)}}^{\max}=\prod_{i_{m}^{\left(\ell\right)}\in S_{m}^{\left(\ell\right)}}\rho_{u,i_{m}^{\left(\ell\right)}}T_{u}^{\rm{CP}}+\left(1-\prod_{i_{m}^{\left(\ell\right)}\in S_{m}^{\left(\ell\right)}}\rho_{u,i_{m}^{\left(\ell\right)}}\right)T_{u,S_{m}^{\left(\ell\right)}}^{\mathrm{out},\max},$ (16)
which denotes the maximal per-token inference latency when user $u$ requests the expert(s) $S_{m}^{\left(\ell\right)}$. The term $T_{u,S_{m}^{\left(\ell\right)}}^{\mathrm{out,max}}$ represents the worst-case latency when the required experts that are not cached locally are also unavailable at any edge server, and thus the token must be processed at the cloud. $T_{u,S_{m}^{\left(\ell\right)}}^{\mathrm{out,max}}$ can be expressed as
$\begin{split}T_{u,S_{m}^{\left(\ell\right)}}^{\mathrm{out,max}}&=T_{u,n_{u},m}^{\mathrm{UL}}+T_{n_{u},{\mathrm{C}},m}+T_{\mathrm{C},m}^{\mathrm{CP}}\\ &+\left(K_{m}-\sum_{i_{m}^{\left(\ell\right)}\in S_{m}^{\left(\ell\right)}}\rho_{u,i_{m}^{\left(\ell\right)}}\right)\left(T_{u,n_{u},m}^{\mathrm{DL}}+T_{{\mathrm{C}},n_{u},m}\right).\end{split}$ (17)
To optimize the expert caching for efficient MoE inference, we aim to minimize the average latency of all generated tokens. Therefore, the objective function $F\left(\mathbf{X}\right)$ can be expressed as
$\begin{split}F\left(\mathbf{X}\right)=\frac{1}{U}\sum_{u\in\mathcal{U}}\sum_{m\in\mathcal{M}}p_{u,m}\sum_{\ell\in\mathcal{L}_{m}}\sum_{S_{m}^{(\ell)}\in{\mathcal{A}}_{m}^{(\ell)}}p_{u,S_{m}^{(\ell)}}r_{u,S_{m}^{(\ell)}}\left(\mathbf{X}\right),\end{split}$ (18)
and the optimization problem can be formulated as follows
$\mathcal{P}1:\;\max_{\mathbf{X}}\;F\left(\mathbf{X}\right)$ (19a)
${\mathrm{s.t.}}\;\sum_{m\in\mathcal{M}}\sum_{\ell\in\mathcal{L}_{m}}\sum_{i_{m}^{\left(\ell\right)}\in\mathcal{E}_{m}^{\left(\ell\right)}}x_{n,i_{m}^{\left(\ell\right)}}b_{m}\leq Q_{n},\;\forall n\in\mathcal{N},$ (19b)
$x_{n,i_{m}^{\left(\ell\right)}}\in\left\{0,1\right\},\forall n\in\mathcal{N},m\in\mathcal{M},\ell\in\mathcal{L}_{m},i_{m}^{(\ell)}\in\mathcal{E}_{m}^{(\ell)},$ (19c)
where Constraint (19b) is a knapsack constraint denoting that the cached experts within any edge server cannot exceed its storage capacity, and Constraint (19c) denotes that the variable $\mathbf{X}=\left\{x_{n,i_{m}^{\left(\ell\right)}}\mid\forall n\in\mathcal{N},i_{m}^{(\ell)}\in\mathcal{E}_{m}^{(\ell)}\right\}$ is a set of binary variables.
IV-B Problem Characterization
Due to the cooperation of multiple edge servers and the correlation of activated experts, there is a complex coupling relationship within the decision variable, making $\mathcal{P}1$ challenging to solve. In this subsection, we examine the properties of $\mathcal{P}1$ and perform problem mappings, providing a basis for the algorithm design in the next section.
We define the following ground set $V$: $V=\{\underbrace{v_{1}^{1},\ldots,v_{E}^{1}}_{V_{1}},\ldots,\underbrace{v_{1}^{N},\ldots,v_{E}^{N}}_{V_{N}}\}$, where $v_{i}^{n}$ is an abstract element denoting the caching state of expert $i$ at the edge server $n$.
Definition 1 (Submodular Function).
Defined over a finite set $V$, a set function $f:2^{V}\rightarrow\mathbb{R}$ is said to be submodular when for every subset $X,Y\subseteq V$, we have $f\left(X\right)+f\left(Y\right)\geq f\left(X\cup Y\right)+f\left(X\cap Y\right)$. Alternatively, defining $f\left(x|X\right)=f\left(X\cup x\right)-f\left(X\right)$ as the marginal gain obtained by incorporating item $x$ into the set $X\subseteq V$, submodularity can equivalently be characterized by the diminishing returns property: for any $X\subseteq Y\subset V$ and $v\in V\setminus Y$, it holds that $f(v|X)\geq f(v|Y)$.
Definition 2 (Supermodular Function).
A set function $f:2^{V}\rightarrow\mathbb{R}$ is said to be supermodular if the function $-f$ is submodular. That is, for every $X,Y\subseteq V$, it holds that: $f(X)+f(Y)\leq f(X\cup Y)+f(X\cap Y)$. Equivalently, $f$ is supermodular if it satisfies the increasing returns property: for all $X\subseteq Y\subseteq V$ and $v\in V\setminus Y$, $f(v\mid X)\leq f(v\mid Y)$ holds.
Proposition 1.
For Top-$K$ strategy, when $K_{m}=1$ for any $m\in\mathcal{M}$, $\mathcal{P}1$ is a monotone non-decreasing submodular maximization problem with $N$ knapsack constraints.
Proof.
Monotonicity is evident since the addition of a new expert placement cannot decrease the value of the objective function $F\left(\mathbf{X}\right)$.
Then, we show the submodularity or supermodularity of the function. First, we prove the submodularity of $F\left(\mathbf{X}\right)$ when $K_{m}=1$ for all MoE models holds. In this case, the probabilities of a user activating different experts within the same layer are independent. Since the summation of submodular functions preserves submodularity, we only need to demonstrate that the set function $F_{u,S_{m}^{(\ell)}}\left(\mathbf{X}\right)=p_{u,S_{m}^{(\ell)}}r_{u,S_{m}^{(\ell)}}\left(\mathbf{X}\right)$ is submodular for a user $u$ and an expert $S_{m}^{(\ell)}$ of model $m$’s $\ell$-th layer. Consider two caching strategy sets, $\dot{X}$ and $\hat{X}$, where $\dot{X}\subset\hat{X}\subset V$ and $V$ is the ground set which has been defined before. Under the caching strategy $\hat{X}$ and $\dot{X}$, user $u$ transmits its token to edge server $\hat{n}$ and $\dot{n}$, respectively, and ${{T_{{n_{u}},\hat{n},m}}+T_{\hat{n},{n_{u}},m}}\leq{{T_{{n_{u}},\dot{n},m}}+T_{\dot{n},{n_{u}},m}}$ holds, indicating that the round-trip transmission latency between edge server $n_{u}$ and $\hat{n}$ is lower than that between $n_{u}$ and $\dot{n}$. Specifically, when $\hat{n}$ or $\dot{n}$ is identical to $n_{u}$, the round-trip transmission latency between edge server $n_{u}$ and $\hat{n}$ or between $n_{u}$ and $\dot{n}$ is 0. Suppose we add an expert $S_{m}^{(\ell)}\in V\setminus\hat{X}$ to the edge server $n$, we discuss the following two cases:
  1. 1.
    When ${{T_{{n_{u}},n,m}}+T_{n,{n_{u}},m}}>{{T_{{n_{u}},\hat{n},m}}+T_{\hat{n},{n_{u}},m}}$: In this case, the marginal value satisfies of caching strategy $\hat{X}$ is given by $F_{u,S_{m}^{(\ell)}}\left(S_{m}^{(\ell)}\mid\hat{X}\right)=0$. If ${{T_{{n_{u}},\dot{n},m}}+T_{\dot{n},{n_{u}},m}}<{{T_{{n_{u}},n,m}}+T_{n,{n_{u}},m}}$, the marginal value of caching strategy $\dot{X}$ remains zero. If ${{T_{{n_{u}},\dot{n},m}}+T_{\dot{n},{n_{u}},m}}>{{T_{{n_{u}},n,m}}+T_{n,{n_{u}},m}}$, the marginal value of $\dot{X}$ is given by $F_{u,S_{m}^{(\ell)}}\left(S_{m}^{(\ell)}\mid\dot{X}\right)=p_{u,S_{m}^{\left(\ell\right)}}\left({{T_{{n_{u}},\dot{n},m}}+T_{\dot{n},{n_{u}},m}}-{{T_{{n_{u}},n,m}}-T_{n,{n_{u}},m}}\right)>0$.
  2. 2.
    When ${{T_{{n_{u}},n,m}}+T_{n,{n_{u}},m}}<{{T_{{n_{u}},\hat{n},m}}+T_{\hat{n},{n_{u}},m}}$: In this case, the marginal value of strategy $\hat{X}$ is given by $F_{u,S_{m}^{(\ell)}}\left(S_{m}^{(\ell)}\mid\hat{X}\right)=p_{u,S_{m}^{\left(\ell\right)}}\left({{T_{{n_{u}},\hat{n},m}}+T_{\hat{n},{n_{u}},m}}-{{T_{{n_{u}},n,m}}-T_{n,{n_{u}},m}}\right)$. The marginal value of strategy $\dot{X}$ is given by $F_{u,S_{m}^{(\ell)}}\left(S_{m}^{(\ell)}\mid\dot{X}\right)=p_{u,S_{m}^{\left(\ell\right)}}\left({{T_{{n_{u}},\dot{n},m}}+T_{\dot{n},{n_{u}},m}}-{{T_{{n_{u}},n,m}}-T_{n,{n_{u}},m}}\right)$. Thus, the difference in marginal gains between $\hat{X}$ and $\dot{X}$ can be expressed as
    $F_{u,S_{m}^{\left(\ell\right)}}\left(S_{m}^{(\ell)}\mid\hat{X}\right)-F_{u,S_{m}^{\left(\ell\right)}}\left(S_{m}^{(\ell)}\mid\dot{X}\right)$
    $=p_{u,S_{m}^{\left(\ell\right)}}\left({{T_{{n_{u}},\hat{n},m}}+T_{\hat{n},{n_{u}},m}}-{{T_{{n_{u}},\dot{n},m}}-T_{\dot{n},{n_{u}},m}}\right)\leq 0,$
which completes the proof.
Proposition 2.
When $K_{m}>1$ for any $m\in\mathcal{M}$, $\mathcal{P}1$ is a monotone non-decreasing non-submodular non-supermodular maximization problem with $N$ knapsack constraints.
Proof.
It is easy to show that monotonicity holds, which is the same as the case of $K_{m}=1$. For non-submodularity and non-supermodularity, we first take $K_{m}=2$ as an example, and then extend this property to a more general case for any $K_{m}>1$.
We consider two caching strategy sets, $\dot{X}$ and $\hat{X}$, where $\dot{X}\subset\hat{X}\subset V$, and introduce an expert $j_{m}^{(\ell)}$ to the set $V\setminus\hat{X}$. Different from $K_{m}=1$, the activated experts appear in pairs when $K_{m}>1$, and the introduction of any expert will affect the latency when requesting the expert pairs involving this expert.
Fig. 5 shows the caching strategy after introducing a new expert $j_{m}^{(\ell)}$ to the sets $\dot{X}$ and $\hat{X}$. Without loss of generality, we assume that the round-trip transmission latency between edge server $A$ (i.e., $n_{u}$) and edge servers $B$, $C$, and $D$ are in an increasing order. For simplicity, we assume that for model $m$’s $\ell$-th MoE layer, the experts except $i_{m}^{(\ell)}$ and $j_{m}^{(\ell)}$ can only be found at the cloud. Also, only the users covered by the edge server $A$ require $S_{m}^{(\ell)}=\left\{i_{m}^{(\ell)},j_{m}^{(\ell)}\right\}$ for inference. In this case, introducing $j_{m}^{(\ell)}$ will only impact the latency when users covered by $A$ require $S_{m}^{(\ell)}$. Therefore, we only need to discuss the property of $F_{u,S_{m}^{(\ell)}}\left(\mathbf{X}\right)=p_{u,S_{m}^{(\ell)}}r_{u,S_{m}^{(\ell)}}\left(\mathbf{X}\right)$ and then obtain the property of $F(\mathbf{X})$.
(a) $F_{u,S_{m}^{(\ell)}}\left(j_{m}^{(\ell)}\mid\hat{\mathbf{X}}\right)-F_{u,S_{m}^{(\ell)}}\left(j_{m}^{(\ell)}\mid\dot{\mathbf{X}}\right)>0$
(b) $F_{u,S_{m}^{(\ell)}}\left(j_{m}^{(\ell)}\mid\hat{\mathbf{X}}\right)-F_{u,S_{m}^{(\ell)}}\left(j_{m}^{(\ell)}\mid\dot{\mathbf{X}}\right)<0$
Figure 5: Different relationships between the marginal values of $\hat{\mathbf{X}}$ and $\dot{\mathbf{X}}$.
First, we discuss the case in Fig. 5(a). In order to minimize the E2E latency, for the set $\hat{X}$, the token will be transmitted to edge server $\{B,C\}$ and $B$, before and after adding $j_{m}^{(\ell)}$ to edge server $B$, respectively. For the set $\dot{X}$, the token will be transmitted to edge server $\{C,D\}$ and $\{B,D\}$, before and after adding $j_{m}^{(\ell)}$ to edge server $B$, respectively. The marginal values of $\hat{X}$ and $\dot{X}$ are given by $F_{u,S_{m}^{(\ell)}}\left(j_{m}^{(\ell)}\mid\hat{X}\right)=p_{u,S_{m}^{(\ell)}}\left(T_{A,C,m}+T_{C,A,m}-T_{B,A,m}+T_{{\rm{E}},m}^{\rm{CP}}\right)$ and $F_{u,S_{m}^{(\ell)}}\left(j_{m}^{(\ell)}\mid\dot{X}\right)=p_{u,S_{m}^{(\ell)}}(T_{A,C,m}+T_{C,A,m}-T_{A,B,m}-T_{B,A,m})$, respectively. In this case, we have $F_{u,S_{m}^{(\ell)}}\left(j_{m}^{(\ell)}\mid\hat{X}\right)-F_{u,S_{m}^{(\ell)}}\left(j_{m}^{(\ell)}\mid\dot{X}\right)>0$, and then $F\left(j_{m}^{(\ell)}\mid\hat{X}\right)-F\left(j_{m}^{(\ell)}\mid\dot{X}\right)>0$ holds.
Then, we discuss the scenario in Fig. 5(b). For the set $\hat{X}$, before and after adding $j_{m}^{(\ell)}$ to edge server $C$, the token will be processed at edge server $B$ to minimize the E2E latency. Thus, the marginal value of $\hat{X}$ is given by $F_{u,S_{m}^{(\ell)}}\left(j_{m}^{(\ell)}\mid\hat{X}\right)=0$. For the set $\dot{X}$, the token will be transmitted to edge server $\{C,D\}$ and $C$, before and after adding $j_{m}^{(\ell)}$ to edge server $C$, respectively. The marginal values of $\dot{X}$ is given by $F_{u,S_{m}^{(\ell)}}\left(j_{m}^{(\ell)}\mid\dot{X}\right)=p_{u,S_{m}^{(\ell)}}(T_{A,D,m}+T_{D,A,m}-T_{C,A,m}+T_{{\rm{E}},m}^{\rm{CP}})$. Thus, we have $F_{u,S_{m}^{(\ell)}}\left(j_{m}^{(\ell)}\mid\hat{X}\right)-F_{u,S_{m}^{(\ell)}}\left(j_{m}^{(\ell)}\mid\dot{X}\right)<0$, and then $F\left(j_{m}^{(\ell)}\mid\hat{X}\right)-F\left(j_{m}^{(\ell)}\mid\dot{X}\right)<0$ holds. At this point, when $K_{m}=2$, we verify that $F\left(j_{m}^{(\ell)}\mid\hat{X}\right)-F\left(j_{m}^{(\ell)}\mid\dot{X}\right)$ is either positive or negative, leading to the property of non-submodularity and non-supermodularity.
Next, we extend this property to a more general case for $K_{m}>1$. The sketch proof is as follows: the expert $i_{m}^{(\ell)}$ in Fig. 5 is replaced by all the required experts excluding $j_{m}^{(\ell)}$, i.e., the expert group including experts $S_{m}^{(\ell)}\setminus j_{m}^{(\ell)}$. With a similar method to the case $K_{m}=2$, we can prove that $F\left(j_{m}^{(\ell)}\mid\hat{X}\right)-F\left(j_{m}^{(\ell)}\mid\dot{X}\right)$ may also be either positive or negative.
Therefore, when $K_{m}>1$, the non-submodularity and non-supermodularity always hold for $\mathcal{P}1$.
V Proposed Algorithms
Based on the structural analysis of $\mathcal{P}1$ under different values of $K$ in the previous section, we divide the algorithmic discussion into two parts: one for the special case of $K=1$, and the other for the general case of $K\geq 1$. The special case $K=1$ is suitable for the Switch Transformer (ST) family, such as ST-based-X with $X=\{8,16,32,64,128\}$ [15] and Hash Layer [40]. The general case $K\geq 1$ applies to GShard and MoE-LLaVA family [18] with the Top-2 strategy and DBRX [41] with the Top-4 strategy. In both cases, we will develop approximate algorithms with provable guarantees. Note that the algorithms for the general case $K\geq 1$ can obviously achieve an approximation guarantee to the special case $K=1$, albeit with a slightly weaker approximation ratio compared to the algorithm designed for $K=1$.
V-A Special Case: $K=1$
We denote the objective function of $\mathcal{P}1$ as $F_{1}\left(\mathbf{X}\right)$ in the special case. Let ${\mathbf{X}}^{\ast}$ denote the optimal solution of $F_{1}\left(\mathbf{X}\right)$.
V-A1 Algorithm design
Since the optimization problem $\mathcal{P}1$ is a submodular maximization problem with knapsack constraints, we can utilize the greedy-based algorithm to obtain the solutions, as shown in Algorithm 1.
Algorithm 1 Greedy-based Algorithm When $K=1$
1:Input: Ground set $V$, objective function $F_{1}$, model $m$’s expert data size $b_{m}$, number of edge servers $N$, storage capacities $Q_{n}$ for $n\in\mathcal{N}$.
2:Output: Expert caching results $S$ (and the way it is cached).
3:Let $S=\emptyset$ and $S=\left\{S_{1},\ldots,S_{N}\right\}$ be its partition into $N$ edge servers.
4:while $V\setminus S\neq\emptyset$ and there exists $n\in\mathcal{N}$ such that $\mathop{\sum}\limits_{i\in S_{n}}b_{i}\leq Q_{n}$ do
5:     $i^{*}=\arg\max_{i\in V\setminus S}\frac{F_{1}(i\mid S)}{b_{i}}$.
6:     $S=S+i^{*}$ and $S_{n}=S_{n}+i^{*}$.
7:end while
8:return $S=\left\{S_{1},\ldots,S_{N}\right\}$.
Proposition 3.
Algorithm 1 returns a expert caching set $S$ that satisfies $F_{1}\left(S\right)\geq\left(1-1/e\right)F_{1}\left({\mathbf{X}}^{\ast}\right)$.
Proof.
The proof follows the approach in [42]. ∎
V-A2 Computational complexity
Since Algorithm 1 is developed from greedy Algorithm, its computational complexity is given by 1 is $\mathcal{O}(N^{2}E^{2})$.
V-B General Case: $K\geq 1$
In this subsection, we extend our discussion to the general case where both $K=1$ and $K>1$ strategies are involved. Recall that for $K>1$, the objective function of $\mathcal{P}1$ is neither submodular nor supermodular, making it challenging to solve. To address this, we first decompose the original problem $\mathcal{P}1$ into $N$ subproblems and solve them successively in the order of edge server index. Based on the analysis of each subproblem, we demonstrate why the greedy algorithm—effective in the special case—fails to provide a constant approximation ratio in the general case. Leveraging the structure of each subproblem, we then propose a dynamic programming (DP)-based algorithm and also an accelerated algorithm to obtain solutions with a constant approximation guarantee. A detailed discussion of the proposed algorithms and their performance is also provided.
V-B1 Problem decomposition
The subproblem of edge server $n$ can be expressed as follows.
$\mathcal{P}2_{n}:\;\max_{\tilde{\mathbf{X}}_{n}}\;{\tilde{F}}_{n}\left(\tilde{\mathbf{X}}_{n}\right)$ (20a)
${\mathrm{s.t.}}\;\sum_{m\in\mathcal{M}}\sum_{\ell\in\mathcal{L}_{m}}\sum_{i\in\mathcal{E}_{m}}x_{n,i_{m}^{\left(\ell\right)}}b_{m}\leq Q_{n},$ (20b)
$x_{n,i_{m}^{\left(\ell\right)}}\in\left\{0,1\right\},\;\forall m\in\mathcal{M},\ell\in\mathcal{L}_{m},i_{m}^{(\ell)}\in\mathcal{E}_{m}^{(\ell)},$ (20c)
The objective function ${\tilde{F}}_{n}\left(\tilde{\mathbf{X}}_{n}\right)$ is given by
${\tilde{F}}_{n}\left(\tilde{\mathbf{X}}_{n}\right)=\sum_{u\in\mathcal{U}}\sum_{m\in\mathcal{M}}p_{u,m}\sum_{\ell\in\mathcal{L}_{m}}\sum_{S_{m}^{(\ell)}\in{\mathcal{A}}_{m}^{(\ell)}}p_{u,S_{m}^{(\ell)}}{\tilde{r}}_{u,S_{m}^{(\ell)}}\left(\tilde{\mathbf{X}}_{n}\right),$ (21)
where
That is, given the caching results of the previous edge servers from $n=1$ to $n-1$, we aim to optimize the caching strategy of edge server $n$ to maximize the average reduction in latency. Specifically, when $n=1$, ${\tilde{r}}_{u,S_{m}^{(\ell)}}\left(\tilde{\mathbf{X}}_{1}\right)=T_{u,S_{m}^{(\ell)}}^{\max}-T_{u,S_{m}^{(\ell)}}^{\mathrm{token}}\left(\tilde{\mathbf{X}}_{1}\right)$. Therefore, the objective function of $\mathcal{P}1$ follows from
$F(\tilde{\mathbf{X}})=F\left(\bigcup_{n\in\mathcal{N}}\tilde{\mathbf{X}}_{n}\right)=\sum_{n\in\mathcal{N}}\tilde{F}_{n}\left(\tilde{\mathbf{X}}_{n}\right).$ (23)
The second equality holds since the sum of $\tilde{F}_{n}$ is a telescoping sum, where all intermediate terms cancel out, leaving only the first and the last terms. This motivates us to divide the original problem into $N$ subproblems and solve them successively.
V-B2 Subproblem analysis for $K\geq 1$
Take the subproblem $\mathcal{P}2_{n}$ as an example, we analyze its property and propose an efficient accelerated knapsack algorithm to solve it.
Proposition 4.
$\mathcal{P}2_{n}$ is maximizing the sum of a non-decreasing modular function and a non-decreasing supermodular function subject to a knapsack constraint. The modular function corresponds to the term including $m$ with $K_{m}=1$ and the supermodular function corresponds to the term including $m$ with $K_{m}>1$.
Proof.
The proof of Proposition 4 is omitted here because the procedure is similar to that of Proposition 2. ∎
First, we provide an explanation about why the traditional greedy algorithm (Algorithm 1) cannot achieve a constant-approximation guarantee when solving $\mathcal{P}2_{n}$.
Remark 1.
(Greedy Limitations in Supermodular Optimization Problems) In MoE inference with $K>1$, each token must activate multiple experts simultaneously within a layer, introducing strong interdependencies among the selected experts. The overall utility of a selection can only be accurately assessed once all required experts have been determined. Prior to this, the contribution of any individual expert is difficult to evaluate in isolation, often resulting in an underestimation of its true utility. While the greedy algorithm may be effective when $K=1$, where expert activations are independent and the utility of each expert can be evaluated individually, they fail to account for the combinatorial nature of utility when $K>1$. In these cases, early greedy selections are made with incomplete information and may restrict the effectiveness of subsequent choices, leading to suboptimal overall performance due to cascading effects. In contrast, the proposed method precomputes a utility score for each expert based on its isolated contribution when added to an empty set. Although the utilities are computed independently, the dynamic programming approach systematically explores the combinatorial space of expert subsets, enabling globally optimal selection. This strategy mitigates the shortsightedness inherent in greedy methods and yields higher-quality solutions.
V-B3 The proposed DP-based algorithm
Since greedy algorithm is improper to solve $\mathcal{P}2_{n}$, we reformulate $\mathcal{P}2_{n}$ into a tractable modular optimization problem and provide solutions.
Let $\tilde{F}_{n}\left(\tilde{\mathbf{X}}_{n}\right)\triangleq{\tilde{F}}_{n}^{\mathrm{mod}}\left(\tilde{\mathbf{X}}_{n}\right)+{\tilde{F}}_{n}^{\mathrm{super}}\left(\tilde{\mathbf{X}}_{n}\right)$, where ${\tilde{F}}_{n}^{\mathrm{mod}}$ and ${\tilde{F}}_{n}^{\mathrm{super}}$ denote the modular and supermodular components of ${\tilde{F}}_{n}$, respectively. We define a modular function as ${H_{n}}:2^{\mathcal{N}}\to\mathbb{R}_{+}$ with $H_{n}\left(\tilde{\mathbf{X}}_{n}\right)=\sum_{z\in\tilde{\mathbf{X}}_{n}}{\tilde{F}}_{n}^{\mathrm{super}}(z\mid\emptyset)$. Then, we reformulate $\mathcal{P}2_{n}$ as $\mathcal{P}3_{n}$, which can be expressed as follows,
$\mathcal{P}3_{n}:\mathop{\max}\limits_{\tilde{\mathbf{X}}_{n}}\;{\bar{F}}_{n}\left(\tilde{\mathbf{X}}_{n}\right)={\tilde{F}}_{n}^{\mathrm{mod}}\left(\tilde{\mathbf{X}}_{n}\right)+H_{n}\left(\tilde{\mathbf{X}}_{n}\right)$ (24a)
[…]
Since the objective function ${\bar{F}}_{n}$ is a modular function, $\mathcal{P}3_{n}$ is a general knapsack problem, which can be solved optimally with a DP-based algorithm in polynomial time. We outline the procedure in Algorithm 2. The gap between the proposed solutions and the global optimal solutions of $\mathcal{P}1$ will be discussed in the following paragraphs.
Algorithm 2 Proposed DP-based Algorithm When $K\geq 1$
1:Expert values ${\tilde{F}}_{n}^{\mathrm{mod}}(i\mid\emptyset)$ or ${\tilde{F}}_{n}^{\mathrm{super}}(i\mid\emptyset)$ for $i\in V_{n}$, expert data size $b_{m(i)}$ for $i\in V_{n}$, edge server $n$’s storage capacity $Q_{n}$, $V_{n}$’s cardinality $E$.
2:Expert caching results $S$ (and the way it is cached).
3:for $n\in\mathcal{N}$ do
4:     Initialize ${\rm{dp}}[0\ldots E][0\ldots Q_{n}]\leftarrow 0$
5:     for $i\in[1,E]$ do
6:         for $q\in[0,Q_{n}]$ do
7:              if $b_{m(i)}>q$ then
8:                  ${\rm{dp}}[i][q]\leftarrow{\rm{dp}}[i-1][q]$
9:              else
10:                  if $K_{m(i)}=1$ then
11:                       ${\rm{dp}}[i][q]\leftarrow\max({\rm{dp}}[i-1][q],\ {\rm{dp}}[i-1][q-b_{m(i)}]+{\tilde{F}}_{n}^{\mathrm{mod}}(i\mid\emptyset))$
12:                  else
13:                       ${\rm{dp}}[i][q]\leftarrow\max({\rm{dp}}[i-1][q],\ {\rm{dp}}[i-1][q-b_{m(i)}]+{\tilde{F}}_{n}^{\mathrm{super}}(i\mid\emptyset))$
14:                  end if
15:              end if
16:         end for
17:     end for
18:     return ${\rm{dp}}[E][Q_{n}]$
19:     // Traceback to get selected experts of edge server $n$
20:     Initialize $\ddot{\mathbf{X}}_{n}\leftarrow\emptyset$, $i\leftarrow E$, $q\leftarrow Q_{n}$
21:     while $i>0$ do
22:         if ${\rm dp}[i][q]\neq{\rm dp}[i-1][q]$ then
23:              $\ddot{\mathbf{X}}_{n}\leftarrow\ddot{\mathbf{X}}_{n}\cup\{i\}$, $q\leftarrow q-b_{m(i)}$
24:         end if
25:         $i\leftarrow i-1$
26:     end while
27:     return $\ddot{\mathbf{X}}_{n}$
28:end for
29:return $\ddot{\mathbf{X}}=\left\{\ddot{\mathbf{X}}_{1},\ldots,\ddot{\mathbf{X}}_{N}\right\}$
V-B4 The proposed accelerated algorithm
Let $T$ denote the distinct data size of experts among all the MoE models. We notice that $T\leq M\ll E$. Here, $\leq$ lies in that numerous downstream models are developed by fine-tuning a pre-trained base model, and $\ll$ comes from the same data size of experts in each MoE model. Based on Definition 3, we design Algorithm 3 with lower complexity than the traditional DP algorithm.
Definition 3.
Let $y^{(i)}$ represent the subsequence of $a$ formed by selecting elements at positions congruent to $i\mod k$, that is, $y^{(i)}=(a_{i},a_{i+k},a_{i+2k},\dots)$. Let $z$ represent a $k$-step concave sequence $z=(b_{0},b_{k},b_{2k},\ldots)$. For any $q$, we can compute the (max,+) convolution of $y^{(i)}$ and $z$ as follows:
$\left(y^{(i)}\oplus z\right)_{q}=\max_{l=0}^{\infty}\{y^{(i)}_{q-l}+z_{l}\}=\max_{l=0}^{\infty}\left\{a_{i+qk-lk}+b_{lk}\right\}.$
Algorithm 3 Proposed Accelerated Algorithm When $K\geq 1$
1:Input: Experts with data size in $\{\tilde{b}_{1},\ldots,\tilde{b}_{T}\}$, expert values ${\tilde{F}}_{n}^{\mathrm{mod}}(i\mid\emptyset)$ or ${\tilde{F}}_{n}^{\mathrm{super}}(i\mid\emptyset)$ for $i\in V_{n}$, storage capacity $Q_{n}$
2:for $n\in\mathcal{N}$ do
3:     Partition items into sets $E_{1},\ldots,E_{T}$, such that $E_{i}=\{j\mid b_{j}={\tilde{b}}_{i}\}$
4:     for $i\in\{1,\ldots,T\}$ and $q\in\{1,\ldots,Q_{n}\}$ do
5:         $s_{q}^{(i)}\leftarrow$ solution for $E_{i}$ with knapsack size $q$
6:     end for
7:     Initialize $\ddot{\mathbf{X}}_{n}$ as an empty sequence
8:     for $i\in\{1,\ldots,T\}$ do
9:         $\ddot{\mathbf{X}}_{n}\leftarrow\ddot{\mathbf{X}}_{n}\oplus s^{(i)}$ using Definition 3
10:     end for
11:     Update $\ddot{\mathbf{X}}_{n}$ by only keeping its first $Q_{n}$ elements
12:end for
13:return $\ddot{\mathbf{X}}=\left\{\ddot{\mathbf{X}}_{1},\ldots,\ddot{\mathbf{X}}_{N}\right\}$
To improve efficiency without sacrificing optimality, Algorithm 3 leverages the observation that the experts can be naturally grouped by their data sizes into a small number of categories. Each subset induces a subsequence whose contributions to the overall solution can be computed independently via efficient convolution operations, as defined in Definition 3. These $T$ (satisfying $T\ll E$) subsequences preserve the structure required by the original DP formulation, allowing their partial solutions to be combined without loss of optimality.
V-B5 Computational complexity
The computational complexity of Algorithm 2 is $\mathcal{O}(NE\left\lceil Q_{n}/(b_{m})_{\min}\right\rceil)$, while the computational complexity of Algorithm 3 $\mathcal{O}(NT\left\lceil Q_{n}/(b_{m})_{\min}\right\rceil))$. The detailed proof procedure is referred to [43].
Compared to the computational complexity of Algorithm 1, the computational of our proposed algorithm is lower since $T$ is much lower than $E$ in MoE models and $\left\lceil Q_{n}/(b_{m})_{\min}\right\rceil)$ is also lower than $E$, especially in the storage-constrained edge networks.
V-B6 Approximation guarantee
In this part, we will provide the approximation guarantee obtained by Algorithm 3 when solving each subproblem $\mathcal{P}2_{n}$ and then give the approximation guarantee of the global solution.
First, the definition of supermodular curvature is provided, followed by the property of supermodular function.
Definition 4.
(Supermodular Curvature [44]) The supermodular curvature of a monotone non-decreasing non-negative supermodular function $g:2^{V}\rightarrow\mathbb{R}$ is defined as $\kappa_{g}=1-\min_{z\in V}\frac{g(z\mid\emptyset)}{g(z\mid V\setminus z)}$.
Supermodular curvature $\kappa_{g}$ is computationally feasible and requires only linear time in the oracle model. Furthermore, we have the following conclusions.
Lemma 1.
(Property of the supermodular function [45, Lemma 2]) If $g:2^{V}\rightarrow\mathbb{R}$ is a non-decreasing supermodular function with curvature $\kappa_{g}$, then for every $A\subseteq V$, it holds that $(1-\kappa_{g})g\left(A\mid\emptyset\right)\leq\sum_{z\in A}g(z\mid\emptyset)\leq g\left(A\mid\emptyset\right)$.
Proposition 5.
Let $\ddot{\mathbf{X}}_{n}\subseteq V$ denote the optimal solution of $\mathcal{P}3_{n}$ obtained by Algorithm 3. We have the following approximation ratio,
${\tilde{F}}_{n}\left(\ddot{\mathbf{X}}_{n}\right)\geq\left(1-(\kappa_{g})_{n}\right){\tilde{F}}_{n}\left(\tilde{\mathbf{X}}_{n}^{\ast}\right),$ (25)
[…]
Proof.
Since the optimal solution of $\mathcal{P}3_{n}$ can be achieved with DP algorithm, ${\bar{F}}_{n}\left(\ddot{\mathbf{X}}_{n}\right)\geq{\bar{F}}_{n}\left(\tilde{\mathbf{X}}_{n}^{\ast}\right)$ holds. According to Lemma 1, we have
$\begin{split}&{\tilde{F}}_{n}^{\mathrm{mod}}\left(\tilde{\mathbf{X}}_{n}\right)+{\tilde{F}}_{n}^{\mathrm{super}}\left(\tilde{\mathbf{X}}_{n}\mid\emptyset\right)\geq{\bar{F}}_{n}\left(\tilde{\mathbf{X}}_{n}\right)\\ &\geq{\tilde{F}}_{n}^{\mathrm{mod}}\left(\tilde{\mathbf{X}}_{n}\right)+\left(1-\kappa_{g}\right){\tilde{F}}_{n}^{\mathrm{super}}\left(\tilde{\mathbf{X}}_{n}\mid\emptyset\right).\end{split}$ (26)
Then, the following inequalities hold,
${\tilde{F}}_{n}\left(\ddot{\mathbf{X}}_{n}\right)\geq{\bar{F}}_{n}\left(\ddot{\mathbf{X}}_{n}\right)+{\tilde{F}}_{n}^{\mathrm{super}}\left(\emptyset\right)$
$\geq{\bar{F}}_{n}\left(\tilde{\mathbf{X}}_{n}^{\ast}\right)+{\tilde{F}}_{n}^{\mathrm{super}}\left(\emptyset\right)$
$\geq{\tilde{F}}_{n}^{\mathrm{mod}}\left(\tilde{\mathbf{X}}_{n}^{\ast}\right)+\left(1-(\kappa_{g})_{n}\right){\tilde{F}}_{n}^{\mathrm{super}}\left(\tilde{\mathbf{X}}_{n}^{\ast}\mid\emptyset\right)+{\tilde{F}}_{n}^{\mathrm{super}}\left(\emptyset\right)$
$={\tilde{F}}_{n}^{\mathrm{mod}}\left(\tilde{\mathbf{X}}_{n}^{\ast}\right)+\left(1-(\kappa_{g})_{n}\right)\left({\tilde{F}}_{n}^{\mathrm{super}}\left(\tilde{\mathbf{X}}_{n}^{\ast}\right)-{\tilde{F}}_{n}^{\mathrm{super}}\left(\emptyset\right)\right)$
$+{\tilde{F}}_{n}^{\mathrm{super}}\left(\emptyset\right)$
$\geq{\tilde{F}}_{n}^{\mathrm{mod}}\left(\tilde{\mathbf{X}}_{n}^{\ast}\right)+\left(1-(\kappa_{g})_{n}\right){\tilde{F}}_{n}^{\mathrm{super}}\left(\tilde{\mathbf{X}}_{n}^{\ast}\right)$
$\geq\left(1-(\kappa_{g})_{n}\right){\tilde{F}}_{n}\left(\tilde{\mathbf{X}}_{n}^{\ast}\right).$
Finally, we establish the following theorem to provide a constant approximation guarantee of our proposed algorithm.
Theorem 1.
Let $\ddot{\mathbf{X}}$ and $\tilde{\mathbf{X}}^{\ast}$ denote the solution obtained by the proposed algorithm for $K\geq 1$ and the global optimal solution of $\mathcal{P}1$. We have $F\left(\ddot{\mathbf{X}}\right)\geq\frac{1-\kappa_{g}^{\max}}{2}F\left(\tilde{\mathbf{X}}^{\ast}\right)$, where $\kappa_{g}^{\max}=\max_{n\in\mathcal{N}}\;(\kappa_{g})_{n}$. Specifically, when $N=1$, the obtained solutions by the proposed algorithm can provide $(1-(\kappa_{g})_{1})$-approximation guarantee.
Proof.
Let $\ddot{\mathbf{X}}_{n}$ and $\tilde{\mathbf{X}}^{\ast}_{n}$ denote the solution obtained by the proposed expert caching algorithm for $K\geq 1$ and the global optimal solution when solving $\mathcal{P}2_{n}$. We introduce a comparison set $\dot{\mathbf{X}}_{n}$, which satisfies
$\dot{\mathbf{X}}_{n}=\left\{\begin{array}[]{l}\tilde{\mathbf{X}}^{\ast}_{n},\quad{\rm{if}}\;\ddot{\mathbf{X}}_{n}\neq\tilde{\mathbf{X}}^{\ast}_{n},\\ \emptyset,\quad{\rm{otherwise}}.\end{array}\right.$ (27)
With (25), we have $\left(1-\kappa_{g}^{\max}\right){\tilde{F}}_{n}\left(\dot{\mathbf{X}}_{n}\right)\leq\left(1-(\kappa_{g})_{n}\right){\tilde{F}}_{n}\left(\dot{\mathbf{X}}_{n}\right)\leq{\tilde{F}}_{n}\left(\ddot{\mathbf{X}}_{n}\right)$ for $\forall n\in\mathcal{N}$. Also, $\left(1-\kappa_{g}^{\max}\right){\tilde{F}}_{n}\left(\tilde{\mathbf{X}}_{n}^{\ast}\right)\leq{\tilde{F}}_{n}\left(\ddot{\mathbf{X}}_{n}\right)+\left(1-\kappa_{g}^{\max}\right){\tilde{F}}_{n}\left(\dot{\mathbf{X}}_{n}\right)$ holds. Therefore, the following inequality holds,
$\begin{split}&\sum_{n\in\mathcal{N}}\left(1-\kappa_{g}^{\max}\right)\tilde{F}_{n}({\tilde{\mathbf{X}}}_{n}^{\ast})\\ &\leq\sum_{n\in\mathcal{N}}{\tilde{F}}_{n}\left(\ddot{\mathbf{X}}_{n}\right)+\left(1-\kappa_{g}^{\max}\right){\tilde{F}}_{n}\left(\dot{\mathbf{X}}_{n}\right)\leq 2\sum_{n\in\mathcal{N}}{\tilde{F}}_{n}\left(\ddot{\mathbf{X}}_{n}\right).\end{split}$ (28)
Then, we have $\sum_{n\in\mathcal{N}}{\tilde{F}}_{n}\left(\ddot{\mathbf{X}}_{n}\right)\geq\frac{1-\kappa_{g}^{\max}}{2}\sum_{n\in\mathcal{N}}\tilde{F}_{n}({\tilde{\mathbf{X}}}_{n}^{\ast})$. Since (23) holds, we have $F\left(\ddot{\mathbf{X}}\right)\geq\frac{1-\kappa_{g}^{\max}}{2}F\left(\tilde{\mathbf{X}}^{\ast}\right)$.
In the special case $N=1$, the approximation ratio can be obtained directly according to Proposition 5. ∎
V-C Discussion of Proposed Algorithm and Solutions
First, we discuss the approximation guarantee in Theorem 1, which affects the gap between the obtained solutions and the theoretical optimum.
Remark 2.
(Applicability of Algorithm 2 and 3 in Special Model Storage Cases) While Algorithm 2 and 3 are designed for $K\geq 1$, they can also be applied to the special case $K_{m}=1$ for any $m\in\mathcal{M}$. If we use Algorithm 2 or 3 to successively solve $\mathcal{P}2_{n}$, the obtained solutions can provide $1/2$-approximation guarantee. The reason lies in that only the modular part is left when solving $\mathcal{P}2_{n}$ for any $n\in\mathcal{N}$, and the optimal solution can be obtained using Algorithm 2 or 3.
Remark 3.
(Estimation of the Approximation Ratio) According to Definition 4, the parameter $\kappa_{g}^{\max}$ in Theorem 1 can be calculated by $\kappa_{g}^{\max}=\frac{T_{n,n^{\prime},\hat{m}}-T_{\mathrm{E},\hat{m}}^{\mathrm{CP}}}{2T_{n,n^{\prime},\hat{m}}-T_{\mathrm{E},\hat{m}}^{\mathrm{CP}}}$ with $\hat{m}=\mathop{\arg\max}\limits_{m\in\mathcal{M}}\;T_{\mathrm{E},m}^{\mathrm{CP}}$ and $n,n^{\prime}=\mathop{\arg\min}\limits_{\begin{subarray}{c}n_{1},n_{2}\in\mathcal{N}\\ n_{1}\neq n_{2}\end{subarray}}\;d_{n_{1},n_{2}}$. In the scenario where communication latency becomes the main bottleneck of total inference latency, i.e., $T_{\mathrm{E},m}^{\mathrm{CP}}$ is much lower than $T_{n,n^{\prime},\hat{m}}$, we have $\kappa_{g}^{\max}\approx\frac{1}{2}$ and $F\left(\ddot{\mathbf{X}}\right)\gtrsim\frac{1}{4}F\left(\tilde{\mathbf{X}}^{\ast}\right)$.
VI Experimental Results
In this section, we conduct simulations to demonstrate the effectiveness of the proposed algorithms.
VI-A Experimental Settings
In our simulations, there are $N=4$ edge servers and $U=20$ edge devices distributed in a cell of size 1 km $\times$ 1 km. The wireless parameters are set as follows: The transmit power of edge device and edge server is 0.01 W [46] and 38 dBm [47], respectively. The bandwidth between edge device and edge server is 5 MHz and the bandwidth between edge servers is 100 MHz. The latency between edge servers and cloud is 0.01 s [48]. The path-loss coefficient is $\alpha=4$, the antenna-related factor is 1, and the noise power spectral density is -174 dBm/Hz [37]. The computing capability of edge devices, edge servers, and cloud is 50 TFLOPs, 82.58 TFLOPs, and 312 TFLOPs, respectively. Besides, the storage capacity of edge servers for expert caching is 5 GB. There are 12 MoE models with more than 1600 experts in our considered MoE model library. Half of them are Switch Transformer family with Top-1 strategy, i.e., ST-b-8, ST-b-16, ST-b-32. The others are MoE-LLaVa family with Top-2 strategy, i.e., MoE-LLaVA-StableLM-1.6B-4e, MoE-LLaVA-Qwen-1.8B-4e, MoE-LLaVA-Phi2-2.7B-4e. Each user has a set of 3 to 5 models to request, and each time it requests one model drawn from a Zipf distribution. Given the limited storage capacity of users and the assumption that the non-expert components of the MoE models are stored locally, each user stores 50 experts from the entire experts of the requested models. The MoE models are tested on the SQA [19] and VQA-v2 [49] datasets.
We adopt the following algorithms as benchmarks:
  • Greedy Algorithm: In each step, the algorithm greedily selects the expert and its corresponding edge server that provides the maximum marginal gain in utility, which corresponds to Algorithm 1.
  • Least Frequently Used (LFU) Algorithm: Each edge server caches the most frequently requested experts until reaching the upper limit of its storage capacity.
  • Random Algorithm Each edge server randomly selects experts to cache until reaching the upper limit of its storage capacity.
VI-B Performance Comparison
This subsection evaluates the latency performance of different algorithms by varying the storage capacities of edge servers $Q$, the number of experts locally stored by each user $\rho$, the distribution of user requests for MoE models $p_{u,m}$, the bandwidth between the user and its associated edge server $B_{u}$, and the number of edge servers $N$ and users $U$.
Fig. 6(a) shows the average latency under different edge server storage capacities. It demonstrates that the proposed algorithm achieves the lowest latency across all storage configurations. Specifically, at a storage capacity of 6.25 GB, the proposed method yields an average latency of 0.057 s, representing a reduction of 11.8% and 57.1% compared to the greedy and LFU algorithms, respectively. The advantages of the proposed algorithm over the greedy algorithm have been analyzed in Remark 1. The reasons that the proposed algorithm outperforms LFU scheme are as follows: We consider the cooperative caching relationship between different edge servers when we adopt successive greedy to decompose the original optimization problem, while the edge servers make their caching decisions independently in LFU algorithm. The Random strategy exhibits the poorest performance in terms of average latency across all storage levels.
(a) Latency vs. storage capacity of edge servers.
(b) Latency vs. number of experts locally stored by each user
Figure 6: Latency performance comparison of algorithms under different storage capacities of edge servers and users.
Fig. 6(b) illustrates the impact of varying the number of experts cached at each user on the average inference latency under different algorithms. As expected, increasing user-side caching capacity leads to latency reduction across all methods, due to decreased reliance on retrieving experts from other nodes, such as edge servers or cloud servers. Additionally, the proposed algorithm consistently outperforms all baselines. More importantly, the performance gap between the proposed and greedy algorithms widens as the number of local cached experts increases, from 6.6% to 12.2%. This trend clearly demonstrates that the proposed method is more capable of fully leveraging user-side caching resources to optimize expert placement, enabling low-latency MoE inference. In contrast, LFU and Random strategies show much higher latency, with limited improvement as user storage increases.
Fig. 7(a) shows the average inference latency under varying numbers of models requested by each user. As expected, the average latency increases for all methods as the number of requested models grows from 2 to 10, due to the higher demand on expert retrieval and increased cache contention. The proposed algorithm consistently achieves the lowest latency across all request levels. Compared with greedy, the latency gap exhibits a non-monotonic trend: it initially widens as the number of requested models increases from 2 to 6, but narrows again as the number exceeds 6. This observation reveals an interesting dynamic in caching effectiveness under increasing user demand. When each user requests only a few models, the request patterns across users are highly diverse. Since the greedy algorithm makes local caching decisions based on immediate benefit, it suffers from higher cache miss rates due to a lack of global awareness. When the number of requested models further increases, user requests tend to overlap more. Consequently, greedy algorithm benefits from this overlap because more frequently requested experts are naturally kept in the cache. As a result, its performance converges toward the proposed method under high request loads. Nonetheless, the proposed algorithm maintains the best overall performance across all settings compared to other benchmarks.
(a) Latency vs. number of models requested per end user.
(b) Latency vs. bandwidth between the user and its associated edge server.
Figure 7: Latency performance comparison of algorithms under varying end-user configurations.
Fig. 7(b) illustrates the average inference latency under different bandwidths between users and their associated edge servers. As the bandwidth increases from 0.5 MHz to 10 MHz, the latency of all methods decreases due to reduced communication delay. The proposed algorithm consistently achieves the lowest latency across all bandwidth levels. While Greedy also benefits from increased bandwidth, the relative latency gap between Greedy and the proposed method widens as bandwidth grows. Specifically, the percentage improvement of the proposed method over Greedy increases steadily from 3.9% at 0.5 MHz to 9.5% at 10 MHz, indicating a growing advantage. This trend occurs because, under low bandwidth, communication delay dominates overall latency, making caching strategies less impactful. As bandwidth increases, transmission bottlenecks diminish, and the system becomes more sensitive to caching efficiency and expert placement. The proposed algorithm, with its globally optimized coordination, is better able to exploit high-bandwidth conditions by optimizing expert placement. These results demonstrate that the proposed method not only performs better under constrained bandwidth but also scales more effectively as bandwidth improves.
Fig. 8(a) illustrates the impact of the number of edge servers on average latency. As expected, increasing the number of edge servers reduces latency for all methods by improving model availability and reducing communication distance between edge servers and users. The proposed algorithm consistently achieves the lowest latency. Notably, the gap between the proposed and greedy algorithms is more significant when the number of edge servers is small, and narrows as more servers are deployed. This is because with fewer edge nodes, efficient global coordination becomes more critical—an advantage of the proposed method. As edge resources increase, even simpler strategies like greedy benefit from more cached experts, reducing the performance gap. Nevertheless, the proposed algorithm remains the most effective across all settings.
(a) Latency vs. number of edge servers.
(b) Latency vs. number of users.
Figure 8: Latency performance comparison of algorithms under different numbers of edge nodes.
Fig. 8(b) shows the average latency under varying numbers of users. As the number of users increases from 10 to 50, the latency of all methods shows a slight rise followed by a decline. This non-monotonic trend is mainly due to the trade-off between resource contention and caching efficiency. In the early stages, a growing number of users leads to more diverse model requests. However, as the user count continues to grow, request patterns begin to overlap, making popular experts more common across users. This improves cache efficiency and reduces latency. The proposed method consistently achieves the lowest latency, showing stronger adaptability to dynamic user distributions. It effectively coordinates cache placement among edge nodes, while LFU and Random algorithms exhibit limited awareness of user behavior. The greedy strategy performs moderately but still lags behind the proposed method.
VI-C Algorithm Running Time Comparison
In this subsection, we discuss the impact of several key parameters on the algorithm’s running time, including the storage capacity of edge server $Q$, the number of MoE models $M$ (equivalent to the number of experts $E$), and the number of edge servers $N$ and users $U$.
Fig. 9 compares the algorithm running time of the proposed and the greedy algorithms under varying edge server’s storage capacities (left) and different numbers of MoE models (right). In both cases, the proposed algorithm demonstrates significantly better computational efficiency. Fig. 9(a) shows the algorithm running time under different edge server storage capacities. As storage increases from 1.25 GB to 7.5 GB, the running time of the greedy algorithm rises sharply, showing poor scalability. In contrast, the proposed algorithm maintains a much lower and more stable runtime over the same range. This striking difference is as follows. Greedy explores the solution space through repeated local searches, and as storage capacity increases, the number of candidate caching combinations grows rapidly, leading to substantial computational overhead. The computational complexity of the proposed method is linear in $Q$, allowing it to scale efficiently with storage size. Fig. 9(b) illustrates the impact of the number of MoE models on the algorithm running time, exhibiting a trend similar to that observed in Fig. 9(a). The nonlinear growth of the greedy algorithm aligns with its computational complexity of $\mathcal{O}(N^{2}E^{2})$.
(a) Effect of storage capacity.
(b) Effect of number of MoE models.
Figure 9: Algorithm running time under different storage capacities and MoE models.
Fig. 10 illustrates the impact of network scale on the algorithm running time, represented by the number of edge servers (left) and users (right). Across both subfigures, the proposed algorithm consistently achieves superior scalability compared to the greedy one. In Fig. 10(a), as the number of edge servers increases, the running time of the greedy algorithm grows rapidly, exceeding 380 s at 10 edge servers. This is due to the combinatorial increase in placement possibilities, which significantly enlarges the search space for greedy’s decision-making. In contrast, the proposed method maintains a much lower execution time. In Fig. 10(b), although users are not part of the computational complexity in both algorithms, increasing the number of users still leads to longer running time for both methods. The greedy algorithm’s running time rises sharply to over 120 s at 50 users, while the proposed algorithm remains efficient, with running time under 8 s. This is because more users introduce a larger volume of requests, which enlarges the input space and introduces overhead in utility calculation.
(a) Effect of number of edge servers.
(b) Effect of number of users.
Figure 10: Algorithm running time under different numbers of nodes.
VII Conclusion Remarks
In this paper, we investigate the expert caching problem to improve the efficiency of MoE inference in distributed networks. Given the caching states of users, we design a latency-minimization optimization problem to optimize the caching strategy at the edge servers. We find that the greedy-based algorithm can solve the problem when $K=1$ with a constant approximation guarantee. However, for the general case when $K\geq 1$, the non-submodularity and non-supermodularity of the problem make the greedy algorithm not applicable. To tackle this challenge, we employ a successive greedy method to decompose the original problem into multiple subproblems, which can be solved using a DP-based algorithm. Notice that the data size of expert networks in a MoE model and fine-tuned variants of a common MoE backbone is identical. We further propose an accelerated algorithm to solve the general problem with a constant approximation ratio.
The proposed expert caching scheme for efficient MoE inference introduces a novel direction for designing expert deployment strategies in storage-constrained distributed edge networks. The current SlimCaching framework can be extended to accommodate more complex scenarios, such as the joint optimization of caching variables across both edge devices and edge servers. Furthermore, expert caching can be integrated with expert prefetching and token batching strategies. In this context, expert caching represents a long-term optimization problem, while expert prefetching and token batching are short-term mechanisms that can be coordinated to further enhance overall system performance.